evaluate ∫ sqrt(1+cos2x / 2) dx

evaluate
∫ sqrt(1+cos2x / 2) dx         

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \sqrt{\frac{1 + cos2x}{2}}dx
cos2x = 2cos^{2}x - 1
1 + cos2x = 2cos^{2}x
\frac{1 + cos2x}{2} = cos^{2}x
I = \int \sqrt{cos^{2}x}dx
I = \int cosx.dx
I = sinx + constant